OpenStudy (anonymous):

derivative of (x^2-4x)/(sqrtx) is it: (3x-4)/(2sqrtx) ??

sam (.sam.):

Answer \[\frac{3x ^{\frac{1}{2}}-\frac{4}{\sqrt{x}}}{2}\]

sam (.sam.):

yeah your answer is correct

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

its correct...